07 Question (5 points) 3rd attempt Dart 112 neintel Feedback See HintPart 1 (2 points) Feedback See Hint Johnny Ocean is considering offering a new service that will send surf conditions to surfPart 3 (1 point) & Feedback See Hint What is the total value to executives of consuming 44.00 surf update emails per year? (R
Answer #1

P 1401 44 44 140

1) The inverse demand function for students is given by
Ps = 44.00 - x
If the email update number is 44 then the surplus is given by the area of the triangle
Ps = (44*44)/2 = 968 or $9.68
Students are willing to pay 968 for this service.

For executives the x is equal to 140
PE = 140 - x
Again, we need area of a triangle
(140 * 140) / 2 = 9800 or $98.00

2) The highest price the student will pay is equal to $9.68
( 44 * 44 ) / 2 = 968

3) The executives are having 44 e-mail updates then we will have to calculate the surplus.
This is the area between the two curves.
\tiny \int_{0}^{44} (140-x)-(44-x)dx = \tiny \int_{0}^{44} 96dx

96 * 44 = 4224

The executive has the value of that option equal to $42.24

4) The surplus for the executive for option 44 updates package is $42.24
The executive is willing to pay 9800 for 140 updates package, so he will need to have that surplus.

98.00 - 42.24 = 55.76

Johnny Ocean could charge a maximum of $55.76 for that service to executive.